OpenStudy (anonymous):

I'm new and this is my first question but here it is. The distance between the end of a lamppost's shadow and the top of the lamppost is 22 feet. The angle measure of the shadow is 15°. I got the answer 33.8.

Directrix (directrix):

Welcome to OpenStudy. sin(15 degrees) = x/22 x = 22 * sin(15 degrees) which is not 33.8 Look at the diagram and see if you think I drew it correctly.  @luvmartialarts